1) For this question, assume that 1980 is the base year. Given macroeconomic conditions in the United States over the past thneed explanations.
Answer #1

1. The answer is A "Real GDP and nominal GDP were equal in 1980"

When we say base year, the Nominal GDP become also real GDP.

2. The answer is C " $ 500 Million Dollars"

Value Added = Value of output - Intermediate goods = 800 - 300 = 500

3. The answer is B "0.3"

Labour Force Participation Rate = Total employed + unemployed / Population * 100

Employed = 20 million

Unemployed = 10 Million

Population = 100 Miilion

so Labour force = 20+10 / 100

= 30 / 100 = 0.3

4. The anwer is E ".33"

The umployment rate = Total unemployed / Labour force

= 10 / 30 = .33